If the expression for the field quantity E in free space is given by: E= E_0 Sin (ωt-kz)a_x,where E_o,ω&k are constants.Find the remaining three field quantities. (i.e. D, B & H)
1
Expert's answer
2020-10-12T07:47:21-0400

"E = E_0 \sin (\omega t-kz)"

In EM wave E and H are connected as

"\sqrt{\epsilon \epsilon_0}\, E = \sqrt{\mu\mu_0}\, H"

So, "H = \sqrt{\frac{\epsilon \epsilon_0}{\mu\mu_0}}\, E = \sqrt{\frac{\epsilon \epsilon_0}{\mu\mu_0}}\, E_0 \sin (\omega t - kz)"

"D = \epsilon \epsilon_0E, \; B = \mu \mu_0 H", therefore

"D = \epsilon \epsilon_0 E_0 \sin (\omega t-kz)" and

"B = \sqrt{\epsilon \epsilon_0 \mu \mu_0} \, E_0 \sin (\omega t-kz)".